OpenStudy (anonymous):

how would i find the first term of a geometric series when an=10, r=2, n=4? thank you so much!

OpenStudy (anonymous):

is an the same thing as the last term

OpenStudy (anonymous):

according to my sheet, an is stated as the nth term

OpenStudy (anonymous):

in geometric series, every term is equal to r*previous_term. so if a1=x then a2=x*r a3=x*r*r a4=x*r*r*r so on. n-th is eqaul to a1*r^(n-1). 4-th term 10=a1*2^3 a1=10/8

OpenStudy (anonymous):

if an is a4 then this is how its done: 10=a1(2)^4-1 So a1=10/8

OpenStudy (anonymous):

thank you all so much!